# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v1

@@ Abstract
ABACUS (Atomic-orbital Based Ab-initio Computation at USTC) is an open-source software for first-principles electronic structure calculations and molecular dynamics simulations. It mainly features density functional theory (DFT) and is compatible with both plane-wave basis sets and numerical atomic orbital basis sets. ABACUS serves as a platform that facilitates the integration of various electronic structure methods, such as Kohn-Sham DFT, stochastic DFT, orbital-free DFT, and real-time time-dependent DFT, etc. In addition, with the aid of high-performance computing, ABACUS is designed to perform efficiently and provide massive amounts of first-principles data for generating general-purpose machine learning potentials, such as DPA models. Furthermore, ABACUS serves as an electronic structure platform that interfaces with several AI-assisted algorithms and packages, such as DeePKS-kit, DeePMD, DP-GEN, DeepH, DeePTB, etc.
@@
Figure 1: ABACUS is dedicated to building an algorithm platform and data engine for AI4ES (AI for Electronic Structure). Since partnering with the open-source community DeepModeling in 2021, ABACUS has garnered over 7,000 commits under the protection of Continuous Integration (CI) tests. Its adaptable architecture allows developers to incorporate numerous electronic structure algorithms. The AI-assisted functional correction method DeePKS permits ABACUS to achieve precise functional results at a reduced cost. Owing to high-performance implementation at various devices, ABACUS effectively produces extensive first-principles data across multiple sectors. Combining advanced electronic structure algorithms and AI-assisted pre-trained model, one can transfer the precision of quantum mechanics across scales.
@@ I Introduction
In recent decades, coupled with the development of algorithms and the monumental leap in computational power, Density Functional Theory (DFT) Hohenberg and Kohn (1964); Kohn and Sham (1965) has achieved immense success in materials science, physics, and chemistry, among others. To meet the needs of different application scenarios, a multitude of unique DFT software has emerged Blaha et al. (2020); Zhang et al. (2020a); Romero et al. (2020); Nakata et al. (2020); Kühne et al. (2020); Tancogne-Dejean et al. (2020); Prentice et al. (2020); Sun et al. (2020); Giannozzi et al. (2020); García et al. (2020); Kresse and Joubert (1999); Frisch et al. (2016). For instance, the well-established plane-wave basis with pseudopotential method can accurately handle systems containing hundreds of atoms Kresse and Joubert (1999); Giannozzi et al. (2020). With advancements in high-performance technologies and algorithms, the plane-wave basis set is now capable of addressing systems with thousands or even tens of thousands of atoms Gygi et al. (2006); Watanabe et al. (2017); Lucignano et al. (2019); Gonze et al. (2020); Feng et al. (2024). Smaller basis sets of localized orbitals can effectively handle systems with thousands of atoms with considerably less computational resources and exhibit lower scaling in small systems Kühne et al. (2020); García et al. (2020). Also, linear-scaling algorithms expand the first-principles calculations to larger scales Nakata et al. (2020); Prentice et al. (2020); Tancogne-Dejean et al. (2020). All these DFT software packages are dedicated to extending the applicability of DFT in terms of precision and scale, which has significantly contributed to the current popularity and success of density functional theory.
The ABACUS project was initiated before 2006 and has achieved a series of progress over the nearly past 20 years Chen, Guo, and He (2010a, 2011a); Li et al. (2016); Lin, Ren, and He (2020); Zheng, Ren, and He (2021); Lin, Ren, and He (2021a, b); Qu et al. (2022a, b); Liu and Chen (2022a); Dai, Jin, and He (2022); Liu et al. (2023a); Lin et al. (2024), supporting both plane-wave (PW) basis and numerical atomic orbitals (NAO) basis for first-principles electronic structure and molecular dynamics calculations. The ABACUS team started to cooperate with the open-source community DeepModeling since 2021 aba (2024a), providing a flexible platform for utilizing various electronic structure methods. For example, based on the plane-wave basis set, low-scaling method stochastic DFT (sDFT) for calculations of warm dense matter has been implemented Liu and Chen (2022a); and for numerical atomic orbitals, meta-GGA Liu et al. (2023a) and DFT+U methods have been realized.
The effort to create such an integrated platform is particularly critical today, given the growing intersection of artificial intelligence (AI) and electronic structure algorithms Zhang et al. (2018a); Wang et al. (2018); Chen et al. (2021, 2022a); Kirkpatrick et al. (2021); Jia et al. (2020); Han, Zhang, and Weinan (2019); Chen et al. (2020); Wen et al. (2022). The field of AI is rapidly advancing in algorithm development, thanks to well-built infrastructure such as PyTorch Paszke et al. (2019) and TensorFlow Abadi et al. (2016). In contrast, algorithm research in electronic structure typically spans several years. In the context of a deep integration between AI and electronic structure, this significant discrepancy in development efficiency makes it challenging for electronic structure to keep pace with cutting-edge AI algorithms. To be a competent electronic structure algorithm platform in the AI era, one must fulfill the following criteria: code flexibility that allows for rapid implementation of new algorithms; code engineering to ensure stability under high-frequency development; and open community to attract a large number of researchers interested in electronic structure and AI to make ongoing contributions. A notably successful case is the swift implementation of the AI-assisted algorithm DeePKS Li et al. (2022a) in ABACUS, which will be addressed in Sec. V.6.
In recent years, the rapid development of generative pre-trained models has offered unprecedented opportunities for unifying models with quantum mechanics. Recently, the Deep Potential team has launched open large atomic model project, namely OpenLAM ope (2024a). Owing to informative testing reports, comprehensive functionality, and high-performance heterogeneous implementation, ABACUS effectively meets the needs of OpenLAM for massive and high-precision data production. As depicted in Fig. 1, ABACUS is committed to building an electronic structure algorithm platform and computing engine for the AI4S (AI for Science) era, combining advanced electronic structure algorithms and AI-assisted algorithms to transfer the precision of quantum mechanics across scales.
In this article, we provide a detailed overview of the latest advancements in ABACUS. Sec. III details the theoretical foundations and practical implementation of Kohn-Sham Density Functional Theory (KS-DFT) within ABACUS, including the exchange-correlation functionals, magnetic calculations, self-consistent field iterations, and modifications for surface chemistry, etc. Sec. IV introduces the algorithms based on the plane wave basis employed in ABACUS for KS-DFT, stochastic DFT (sDFT), Orbital-Free DFT (OF-DFT), and its progress in AI-assisted kinetic energy functional. Sec. V discusses methods such as Hybrid Functionals, the DFT+U technique, Real-Time Time-Dependent DFT (RT-TDDFT), and the AI-augmented functional correction method DeePKS, all of which rely on the numerical atomic orbitals basis. Sec. VI concentrates on AI-assisted molecular dynamics and describes ABACUS’s role in generating data for the pre-trained model OpenLAM. Meanwhile, Sec. VII presents the interface between ABACUS and other softwares, while Sec. VI.1 introduces the ongoing pseudopotential orbital library project. Finally, Sec. VIII summarizes the whole review.
II Overview of ABACUS
@@
Figure 2: Code architecture of ABACUS. For users, a few input files need to be prepared in advance, including INPUT, STRU, KPT, pseudopotential files, and orbital files (only for numerical atomic orbitals calculations). For developers who are new to ABACUS, the software features a highly modular design that allows for the swift integration of new functionalities. The core development team is responsible for formulating and refining key data structures, mathematical routines, and other foundational components in ABACUS. For instance, the HContainer module is utilized for storing sparse matrices under the numerical atomic orbital basis set, such as Hamiltonian and density matrices; the Grid Integral module is implemented for grid integration.
Despite ABACUS supporting two types of basis sets and various DFT algorithms, the input file format required for these computations is uniform. A typical ABACUS PW calculation requires the user to prepare a few input files in advance. INPUT specifies the basic parameters of the calculation. Users can select the type of DFT method to enable through the keyword esolver_type, with options including KS-DFT, OF-DFT, sDFT, RT-TDDFT, etc. On the other hand, the keyword calculation determines the type of computation, such as self-consistent field calculation, relaxation, and other options. ABACUS only supports pseudopotential first-principles calculations, thus the pseudopotential files are indispensable. ABACUS supports both norm-conserving and ultrasoft pseudopotentials in various formats.
STRU defines the crystal cell, atomic coordinates, and initial magnetic moments. ABACUS adopts periodic boundary conditions (PBCs) for both crystals and finite systems. For isolated systems, such as atoms, molecules, clusters, etc., one can use a large vacuum layer to avoid interactions from atoms in neighboring cells. In ABACUS, the Ewald method is used to solve the divergence problem caused by PBC Martin (2020).
KPT specifies the k-point sampling of the first Brillouin zone. However, it can be omitted if the user sets the keyword kspacing in INPUT. ABACUS offers three versatile templates for k-point configuration. The first mode reads the number of k-point grids in three dimensions and uniformly samples the Brillouin zone using the Monkhorst-Pack method Monkhorst and Pack (1976), with the option to center k-point sampling at the Gamma point or not. In the second mode, users provide the coordinates and weights of specific k-points directly. The third mode is designed for band structure calculations, where users designate high-symmetry points and the number of k-points between them. For larger systems, ABACUS’s Gamma-Only mode employs real wavefunctions to save time and resources.
ABACUS PW calculations can be activated with the input settings introduced above and optionally with basis_type=pw in the INPUT file. For ABACUS LCAO (Linear combination of atomic orbitals) calculation, alternatively the basis_type=lcao, and additional numerical atomic orbital files *.orb are needed. We provide fully ready-to-use orbital files for users on the official website aba (2024b).
For new feature developers, ABACUS offers highly modular code that allows one to implement new functionalities quickly in a Lego-like assembly way. For core developers, a clear dependency ensures that different modules can be optimized locally. All joint efforts are intended to make ABACUS a user- and developer-friendly software.
III Kohn-Sham Density Functional Theory
III.1 Kohn-Sham Equation
In 1964, Hohenberg and Kohn Hohenberg and Kohn (1964) demonstrated that the energy of a non-degenerate ground state can be uniquely obtained from a density functional of the electronic density of this system, whereas the exact form of such density functional is unknown. In 1965, Kohn and Sham Kohn and Sham (1965) proposed what is later referred to as the Kohn-Sham density functional theory (KS-DFT), where a Slater determinant is used to compute the non-interacting kinetic energy, as well as the densities for computing the electron-nuclear, classical electron-electron, and the non-classical energy. The non-classical energy is obtained from an exchange-correlation functional of density and other ingredients (see discussions in Sec. III for details of the ingredients), so the non-classical energy is more often called the exchange-correlation energy. In KS-DFT, the orbitals are obtained by solving the KS equation
$\left(-\frac{1}{2}\nabla^{2}+\hat{V}_{\rm KS}\right)|\psi_{j}\rangle=\varepsilon_{j}|\psi_{j}\rangle,$ (1)
@@ III.1 Kohn-Sham Equation
where $|\psi_{j}\rangle$ is a Kohn-Sham orbital with energy $\varepsilon_{j}$, and $\hat{V}_{\rm KS}$ is called the KS potential:
$\hat{V}_{\rm KS}=\hat{V}_{\rm ext}+\hat{V}_{\rm H}+\hat{V}_{\rm xc},$ (2)
where the three terms on the right-hand side are called the external-filed term, the Hartree term, and the exchange-correlation (XC) term, respectively. In ABACUS, the Hartree potential is calculated via fast Fourier transforms under periodic boundary conditions Martin (2020).
@@
The KS orbitals in Eq. 1 can be expressed on various types of basis functions, such as plane waves Kresse and Joubert (1999); Giannozzi et al. (2020); Romero et al. (2020), numerical atomic orbitals García et al. (2020); Boker et al. (2011); Nakata et al. (2020); Blum et al. (2009), real-space grids Chelikowsky, Troullier, and Saad (1994); Tancogne-Dejean et al. (2020), Gaussian orbitals Frisch et al. (2016); Kühne et al. (2020), wavelet basis sets Ratcliff et al. (2020), etc. ABACUS supports both plane waves (PWs) and numerical atomic orbitals (NAOs) as basis sets, which are introduced in Sec. IV and Sec. V, respectively.
III.1.1 Pseudopotentials
The external-field term, in absence of external electric field, is computed from a pseudopotential term. First-principles pseudopotentials are essential tools to increase the efficiency in quantum mechanical calculations. A pseudopotential approximates the effect of core electrons and represents the ion core through a simpler potential, which allows valence electrons to be treated explicitly within manageable computational resources.
Several forms of pseudopotentials are now widely used in electronic structure calculations. For example, the norm-conserving pseudopotentials (NCPP) Kleinman (1980); Bachelet and Schlüter (1982) ensure that pseudo-wavefunctions coincide with true wavefunctions outside a certain cutoff radius, while the norm of pseudo-wavefunctions is conserved.
The pseudopotential, $\hat{V}^{\rm PS}$, consists of a local part and a nonlocal part, namely,
$\hat{V}^{\rm PS}=V_{\rm local}(r)+\hat{V}^{\rm KB}.$ (3)
Here, the nonlocal operator is written in the fully nonlocal form proposed by Kleinman and Bylander (KB) Kleinman and Bylander (1982)
$\hat{V}^{\rm KB}$ $=\sum_{ll^{\prime}mm^{\prime}}D_{ll^{\prime}}|\beta^{\rm KB}_{lm}\rangle\langle\beta^{\rm KB}_{l^{\prime}m^{\prime}}|,$ (4)
which projects $D_{ll^{\prime}}$ onto a set of pseudo-atomic orbital bases, $\beta^{\rm KB}_{lm}$, with an angular quantum number $l$ and a magnetic quantum number $m$. The pseudo-atomic-orbital basis function is expressed as a product of a spherical harmonic function and a radial function, namely
$\beta^{\rm KB}_{lm}(\mathbf{r})=\phi^{\rm ps}_{lm}(r)Y_{lm}(\theta,\phi).$ (5)
Furthermore, the ultrasoft pseudopotentials (USPP) Kresse and Hafner (1994); Kresse and Joubert (1999) allow for higher computational efficiency by relaxing the norm-conservation condition. The result is more flexibility in the pseudo-wavefunction within the core region, which in turn means that softer waves can be used. This yields faster-converging basis sets but requires a more sophisticated approach to solve the Kohn-Sham equation.
III.1.2 Exchange-Correlation Functional
Many density functional approximations (DFAs) have been proposed since the publication of KS-DFT. The earliest realization of KS-DFA is local density approximation (LDA), which is the functional of the spin-summed electron density $\rho(\mathbf{r})$. One also finds that using spin density components, namely $\rho_{\uparrow}$ and $\rho_{\downarrow}$, together with a broken-symmetry Slater determinant, gives correct dissociation limit of H2, and the resulting DFA is called local spin density approximation (LSDA).
Density gradients, or spin density gradients, were then introduced as an ingredient to the density functional, resulting in a generalized gradient approximation (GGA) functional. The accuracy of GGAs is greatly improved from the LDA or LSDA functionals, for example, PBEPerdew, Burke, and Ernzerhof (1996) reduces mean absolute error (MAE) of the atomization energies for 20 main-group molecules to 7.9 kcal/mol, whereas the MAE of LSDA is 31.4 kcal/mol.
The next generation of successful DFA is called meta-GGA, to which the introduction of the kinetic energy density, $\tau(\mathbf{r})$, shows improved accuracy compared with GGAs. A widely-used meta-GGA functional is the strongly constrained and appropriately normed (SCAN) functional Sun, Ruzsinszky, and Perdew (2015).
The above-mentioned DFAs use only the local ingredients, which depend solely on a grid point $\mathbf{r}$ in space. However, one may also include a fraction of exact Hartree-Fock exchange energy, and the resulting functional is called hybrid DFA, such as the HSE03 functionalHeyd, Scuseria, and Ernzerhof (2003).
In ABACUS, LDA (and LSDA), GGA, and meta-GGA functionals are implemented with both the plane-wave and the atomic-orbital basis sets. The LDA and GGA functionals are available from both home-built codes and the Libxc library Lib (2024). Meta-GGA functionals, however, can only be used with Libxc in the current version of ABACUS. We refer audiences to Ref. Liu et al., 2023a for testings of SCAN, rSCAN Bartók and Yates (2019), and r2SCAN Furness et al. (2020) on ABACUS. Hybrid functionals are available only with the LCAO basis set in combination with the Libxc and the local resolution of identity (LRI) libraries (More discussions can be found in Sec. V.5 on how hybrid functionals are computed in ABACUS).
III.1.3 Collinear and Non-Collinear Spin
Extending the original spinless KS-DFT to incorporate collinear spin densities involves assigning distinct treatment to the spin-up and spin-down electron populations. In this formalism, each electron is strictly aligned either parallel (spin-up) or antiparallel (spin-down) to a chosen quantization axis. Spin densities for spin up and down channels are defined as $\rho_{\uparrow}(\mathbf{r})$ and $\rho_{\downarrow}(\mathbf{r})$, respectively. The total electron density is
$\rho(\mathbf{r})=\rho_{\uparrow}(\mathbf{r})+\rho_{\downarrow}(\mathbf{r})$ (6)

### v3

@@ Abstract
ABACUS (Atomic-orbital Based Ab-initio Computation at USTC) is an open-source software for first-principles electronic structure calculations and molecular dynamics simulations. It mainly features density functional theory (DFT) and molecular dynamics functions and is compatible with both plane-wave basis sets and numerical atomic orbital basis sets. ABACUS serves as a platform that facilitates the integration of various electronic structure methods, such as Kohn-Sham DFT, stochastic DFT, orbital-free DFT, and real-time time-dependent DFT, etc. In addition, with the aid of high-performance computing, ABACUS is designed to perform efficiently and provide massive amounts of first-principles data for generating general-purpose machine learning potentials, such as DPA models. Furthermore, ABACUS serves as an electronic structure platform that interfaces with several AI-assisted algorithms and packages, such as DeePKS-kit, DeePMD, DP-GEN, DeepH, DeePTB, etc.
@@
Figure 1: ABACUS is dedicated to building an algorithm platform and data engine for AI4ES (AI for Electronic Structure). Since partnering with the open-source community DeepModeling in 2021, ABACUS has garnered over 7,300 commits under the protection of Continuous Integration (CI) tests. Its adaptable architecture allows developers to incorporate numerous electronic structure algorithms. The AI-assisted functional correction method DeePKS permits ABACUS to achieve precise functional results at a reduced cost. Owing to high-performance implementation at various devices, ABACUS effectively produces extensive first-principles data across multiple sectors. Combining advanced electronic structure algorithms and AI-assisted pre-trained models, one can transfer the precision of quantum mechanics across scales.
@@ I Introduction
The theoretical foundations of density-based methodologies can be traced to the Thomas-Fermi (TF) model proposed in 1927. The TF model established the first kinetic energy functional based on the uniform electron gas approximation Thomas (1927); Fermi (1927). Although this orbital-free (OF) framework pioneered the conceptual basis for density functional theory (DFT), its oversimplified treatment of kinetic energy limited chemical accuracy Wang and Carter (2002). In 1964, Hohenberg and Kohn Hohenberg and Kohn (1964) established two theorems. First, the external potential of a many-electron system is uniquely determined by its electron density. Second, a universal energy functional exists whose minimum value corresponds to the ground-state electron density. Later in 1965, Kohn and Sham Kohn and Sham (1965) developed the Kohn-Sham density functional theory (KS-DFT), whose accuracy is determined by the exchange-correlation functional. This framework allows a non-interacting electron system where the kinetic energy can be computed exactly via electronic wavefunctions. Crucially, the non-interacting system can yield the electron density of the real system via the so-called self-consistent field (SCF) method.
Over the past few decades, driven by developments of the DFT method itself and continuous refinement of numerical algorithms, various popular KS-DFT packages emerge Blaha et al. (2020); Zhang et al. (2020a); Romero et al. (2020); Nakata et al. (2020); Kühne et al. (2020); Tancogne-Dejean et al. (2020); Prentice et al. (2020); Sun et al. (2020); Giannozzi et al. (2020); Garc´ıa et al. (2020); Kresse and Joubert (1999); Frisch et al. (2016). In addition, with the exponential growth in computing capabilities, the KS-DFT method has become a cornerstone to predict various properties of materials in physics, chemistry, materials science, biology, geophysics, etc. Despite these achievements, challenges still remain. For example, the community increasingly demands DFT implementations that combine enhanced computational efficiency with improved hardware adaptability across different architectures. Recently, with rapid advancements in artificial intelligence (AI) technologies, an important question arises: can these new technologies drive substantial improvements in computational accuracy and efficiency for electronic structure methods such as KS-DFT and OF-DFT?
In this review paper, we introduce the ABACUS package, which was initiated around 2006 and has achieved a series of progress over the past two decades Chen, Guo, and He (2010a, 2011a); Li et al. (2016); Lin, Ren, and He (2020); Zheng, Ren, and He (2021); Lin, Ren, and He (2021a, b); Qu et al. (2022a, b); Liu and Chen (2022a); Dai, Jin, and He (2022); Liu et al. (2023); Lin et al. (2024). For instance, the ABACUS package supports both plane-wave (PW) basis and numerical atomic orbitals (NAO) basis for first-principles electronic structure calculations, geometry relaxation and molecular dynamics simulations. In addition, ABACUS offers robust support for highly efficient parallelization of diverse electronic structure methods, leveraging MPI, OpenMP, and CUDA to optimize performance across distributed and accelerated computing architectures. Based on recent advancements in AI technologies for atomistic simulations and electronic structure methods Zhang et al. (2018a); Wang et al. (2018); Chen et al. (2021, 2022a); Kirkpatrick et al. (2021); Jia et al. (2020); Han, Zhang, and Weinan (2019); Chen et al. (2020); Wen et al. (2022), researchers have identified significant potential for AI to transform both electronic structure methods and their corresponding software implementations. In this context, the effort to create an integrated platform that integrates physics-based models, high-performance computing, and AI capabilities has become particularly critical. Since 2021, the ABACUS development team has maintained an ongoing collaboration with the DeepModeling open-source community aba (2024).
As illustrated in Fig. 1, ABACUS aims to establish an electronic structure algorithmic platform and computational engine tailored for the AI for Science (AI4S) paradigm. This integration features three aspects. First, state-of-the-art electronic structure method developments with AI-assisted algorithms. Second, molecular dynamics simulations to provide data for the OpenLAM project ope (2024a), which is designed to generate a general-purpose machine learning model for the periodic table. Third, the platform can be potentially used in various industry fields such as battery, catalyst, electronics, ferroelectricity, magnetic materials, high-entropy alloys, etc.
In this work, we provide a detailed overview of the latest developments in ABACUS. Sec. II details implementation of general features in ABACUS, such as the pseudopotentials, exchange-correlation functionals, SCF method, geometry relaxation and molecular dynamics, etc. Sec. III describes the implementation of the KS-DFT, stochastic DFT (sDFT), and OF-DFT methods with the plane wave basis. Moreover, recent advances for AI-assisted kinetic energy density functionals (KEDFs) are introduced for OFDFT. Sec. IV introduces the linear combination of atomic orbitals (LCAO) calculation methods, where numerical atomic orbitals are served as basis to solve the KS equation and yield forces and stress, as well as newly developed methods such as AI-assisted electronic structure method DeePKS, DFT+U, real-time time-dependent DFT (RT-TDDFT) and hybrid functionals, all of which are based on the numerical atomic orbital basis. Sec. V mainly focuses on AI-assisted molecular dynamics methods and describes the roled played by ABACUS. Sec. VI presents various software interfaces between ABACUS and other packages. Sec. V.1 introduces the pseudopotential and numerical atomic orbital library. Finally, we conclude in sec. VII.
@@
Figure 2: Code architecture of ABACUS. For users, a few input files need to be prepared in advance, including INPUT, STRU, KPT, pseudopotential files, and orbital files (only for numerical atomic orbitals calculations). For developers, the software features a highly modular design that allows for the swift integration of new functionalities. The development team is responsible for formulating and refining key data structures, mathematical routines, and other foundational components in ABACUS. For instance, the HContainer module is utilized for storing sparse matrices under the numerical atomic orbital basis set, such as Hamiltonian and density matrices; the Grid Integral module is implemented for grid integration.
II Kohn-Sham Density Functional Theory
II.1 Kohn-Sham Equation
In KS-DFT, the electronic orbitals are obtained by solving the KS equation
$\left[-\frac{1}{2}\nabla^{2}+\hat{V}_{\rm KS}\right]\psi_{i}=\varepsilon_{i}\psi_{i},$ (1)
@@ II.1 Kohn-Sham Equation
where $\psi_{i}$ is the Kohn-Sham wave function with energy $\varepsilon_{i}$, and $\hat{V}_{\rm KS}$ is an effective potential written as
$\hat{V}_{\rm KS}=\hat{V}_{\rm ext}+\hat{V}_{\rm H}+\hat{V}_{\rm xc}.$ (2)
Here the three terms on the right-hand side are the external potential term, the Hartree term, and the exchange-correlation (XC) term, respectively. Typically, the external potential term includes the ion-electron interactions that can be described via pseudopotentials, as well as external fields like the electric field.
@@
The KS equation can be solved via different types of basis sets, such as plane waves Kresse and Joubert (1999); Giannozzi et al. (2020); Romero et al. (2020), numerical atomic orbitals Garc´ıa et al. (2020); Boker et al. (2011); Nakata et al. (2020); Blum et al. (2009), real-space grids Chelikowsky, Troullier, and Saad (1994); Tancogne-Dejean et al. (2020), Gaussian orbitals Frisch et al. (2016); Kühne et al. (2020), wavelet basis sets Ratcliff et al. (2020), etc. Currently, ABACUS supports both plane waves and NAOs as basis sets. On the one hand, the plane wave basis can accurately and efficiently handle condensed matter systems with a relatively small number of atoms. On the other hand, the NAOs are centered on atoms and the required number of basis is substantially less than the plane wave basis. Therefore, the NAO basis is suitable to simulate systems containing thousands of atoms Kühne et al. (2020); Garc´ıa et al. (2020).
II.1.1 Pseudopotentials
The Coulomb potential exhibits a singularity at the atomic nucleus, necessitating an extensive basis set for the precise representation of “near-core” wavefunctions, which imposes substantial computational overhead. To overcome this challenge, several pseudopotential methodologies have been developed in the past few decades Kleinman (1980); Bachelet and Schlüter (1982); Rappe et al. (1990); Vanderbilt (1990); Troullier and Martins (1991); Kresse and Hafner (1994); Kresse and Joubert (1999); Hamann (2013). These pseudopotentials achieve significant reduction in computational costs while preserving computational accuracy.
ABACUS supports both norm-conserving pseudopotentials (NCPPs) Hamann, Schlüter, and Chiang (1979) and ultrasoft pseudopotentials (USPPs) Vanderbilt (1990). Specifically, the pseudopotential operator $\hat{V}_{\rm ps}$ is decomposed into local and nonlocal components in NCPPs as
$\hat{V}_{\rm ps}=\hat{V}_{\rm local}+\hat{V}_{\rm KB},$ (3)
where the local pseudopotential $\hat{V}_{\rm local}$ equals the all-electron functions beyond a cutoff radius. Furthermore, $\hat{V}_{\mathrm{KB}}$ is composed of a set of separable nonlocal projectors as proposed by Kleinman and Bylander Kleinman and Bylander (1982), lated extended by Hamann Hamann (2013). The formula is
$\hat{V}_{\mathrm{KB}}=\sum_{ll^{\prime}mm^{\prime}}D_{ll^{\prime}}|\beta_{lm}\rangle\langle\beta_{l^{\prime}m^{\prime}}|.$ (4)
Here $\beta_{lm}$ is a projector with angular quantum number $l$ and magnetic quantum number $m$, the coefficients are in the $D_{ll^{\prime}}$ matrix. In addition, these projects $\beta_{lm}$ are expressed as the product of a radial function and a spherical harmonic function. In cases such as transition metals or magnetic systems, the non-linear core corrections Louie, Froyen, and Cohen (1982) are usually adopted in the pseudopotential file to improve the transferability of pseudopotentials. For the spin-orbital coupling effect, it can be incorporated into the non-local projectors in the $\hat{V}_{\mathrm{KB}}$ term Theurich and Hill (2001); Corso and Conte (2005); Cuadrado et al. (2021).
Compared to NCPPs, the Ultrasoft pseudopotentials Vanderbilt (1990) offer substantially improved computational efficiency by removing the norm-conservation constraint, which allows greater flexibility in constructing the pseudo-wavefunction within the core region, yielding inherently softer wavefunctions that reduce computational overhead. As a result, USPPs facilitate faster convergence of basis sets expansions, though this advanrage comes with the trade-off of requiring more sophisticated implementation with KS-DFT packages.
II.1.2 Exchange-Correlation Functionals
In KS-DFT codes, to describe collinear spin of given systems, the spin-up electron density $\rho^{\uparrow}(\mathbf{r})$ and spin-down electron density $\rho^{\downarrow}(\mathbf{r})$ are usually treated separately, and then summed up to the total electron density
$\rho(\mathbf{r})=\rho^{\uparrow}(\mathbf{r})+\rho^{\downarrow}(\mathbf{r})$ (5)